A box has 10marbles in it, 7red and 3white. Suppose we draw a marble from the box, replace it, and then draw another. Find the probability that:
(a) Both marbles are red.
(b) Just one of the two marbles is red.

Answer by stanbon(75887)

A box has 10marbles in it, 7red and 3white. Suppose we draw a marble from the box, replace it, and then draw another. Find the probability that:
(a) Both marbles are red.
P(r and r) = 7/10 *7/10 = 49/100 = 0.49
------------------------
(b) Just one of the two marbles is red.
Note: P(w and w) = (3/10)(3/10) = 0.09
-----
P(one is red) = 1 - P(both are red)-P(both are white) = 1-0.49-0.09 = 0.42
--------------------
Cheers,
Stan H.
-------------